# Prompting is Not All You Need! Evaluating LLM Agent Simulation Methodologies with Real-World Online Customer Behavior Data

### v6

@@ Abstract
Recent research shows that LLMs can simulate “believable” human behaviors to power LLM agents via prompt-only methods. In this work, we focus on evaluating LLM’s objective “accuracy” rather than the subjective “believability” in simulating human behavior, leveraging a large-scale, real-world dataset collected from customers’ online shopping actions. We present the first comprehensive evaluation of state-of-the-art LLMs (e.g., DeepSeek-R1, Llama, and Claude) on the task of web shopping action generation. Our results show that out-of-the-box LLM-generated actions are often misaligned with actual human behavior, whereas fine-tuning LLMs on real-world behavioral data substantially improves their ability to generate accurate actions compared to prompt-only methods. Furthermore, incorporating synthesized reasonings into model training leads to additional performance gains, demonstrating the value of explicit rationale in behavior modeling. This work evaluates state-of-the-art LLMs in behavior simulation and provides actionable insights into how real-world action data can enhance the fidelity of LLM agents.
[…]
Prompting is Not All You Need! Evaluating LLM Agent Simulation Methodologies with Real-World Online Customer Behavior Data
@@ 1 Introduction
Recent advances in large language models (LLMs) have enabled the simulation of “believable” human behavior across a range of applications, including web navigation actions Gur et al. (2023); Zhou et al. (2024), social interaction behaviors Park et al. (2023), interpersonal trust behaviors Xie et al. (2024), and user interface interactions Taeb et al. (2024). These developments have sparked growing interest in using LLMs to power agentic systems that emulate human behaviors (LLM Agent) Chen et al. (2025). Despite these successes, current evaluations Park et al. (2023) primarily emphasize subjective measures of “believability” (“how much people feel it is like a human”) instead of the objective “accuracy” (“how much it acts like a human”). The most relevant works that measure the objective model accuracy focus only on the final outcome of a task (e.g., purchasing the final product or not Yao et al. (2023), or ultimately trusting the partner or not Xie et al. (2024)), rather than rigorously evaluating whether the generated sequence of actions mirrors authentic human behavior or not. Consequently, the field currently lacks a robust and quantitative understanding for assessing LLMs at the process-centric, action-level simulation of human behaviors.
@@
In this paper, we focus specifically on the human behavior simulation task, which is to generate the next most likely user action based on the current observation and the history of past actions. For instance, in an online shopping scenario, the model observes the current webpage context (e.g., a product list) and the user’s action history (e.g., previous clicks or queries), and generates the next plausible action a human would take. Real human users often implicitly form reasoning traces behind their action behaviors Cao et al. (2023).
@@
To bridge this gap, our work provides the first systematic evaluation of SOTA LLMs’ accuracy in process-centric, action-level behavior simulation tasks. We leverage a large-scale, real-world dataset consisting of 31,865 user sessions of 3,526 users from an online shopping platform. Each session (Figure 1) comprises a series of timestamp-aligned <context, action> pairs, where the context reflects the webpage observed by the user (e.g., product views, filter states), and the action denotes user inputs such as clicks, searches, or session termination actions. In total, the dataset has 230,965 user actions, and the final outcomes of the sessions include 4,432 purchase actions and 27,433 session termination actions. This dataset enables us to rigorously evaluate how accurately various LLMs can generate human-like behaviors at the action level. We evaluate different models on the next action generation task, benchmarking both the accuracy of generated actions throughout the session and the F1 score for final action generation, following protocols similar to existing work on modeling online shopping behavior.
@@
Beyond evaluation, our dataset uniquely positions us to fine-tune LLMs to enhance their accuracy in behavior simulation tasks. While prior work has primarily relied on prompt-based approaches, we show that model adaptation through fine-tuning leads to significantly better accuracy in action generation and session outcome prediction. Furthermore, drawing inspiration from reasoning-augmented modeling DeepSeek-AI et al. (2025), we hypothesize that exposing models to intermediate reasoning traces–even if synthetically generated–can enhance their ability to simulate human behavior. To test this hypothesis, we augment our dataset with synthesized reasonings from action traces using Claude 3.5 Sonnet and fine-tune models using this augmented data (<context, action, reasoning> triplets) to learn how to generate not only accurate actions but also the underlying reasoning. Our results show that this reasoning-augmented fine-tuning further boosts model performance, highlighting the importance of modeling not just what humans do, but also why they do it.
In summary, this paper has two contributions:
@@
We provide the first quantitative and process-centric evaluation of LLMs to simulate human web action behaviors using real-world data from an online shopping context.
Through our evaluation, we empirically demonstrate that fine-tuning LLMs with synthesized reasoning significantly improves model’s performance, highlighting the value of incorporating explicit reasoning to improve the fidelity of LLM agent applications for role-playing or behavior simulation .
@@ 2.1 Simulation of Human Behavior with LLM
However, the evaluation of these works remains limited in scope. Some focus on the subjective believability of process-centric action traces. For instance, Lu et al. (2025a) conducted qualitative interviews to assess participants’ perceptions of the realism of their UXAgent system. Similarly, Park et al. (2023) proposed an evaluation framework that identified emergent social behaviors among generative agents. On the other hand, works that pursue objective evaluation often do so in a single-shot, outcome-centric manner. Zhou et al. (2024) introduced WebArena, a controlled environment for benchmarking web agents based on task completion rates. ReAct Yao et al. (2023) measured success rates in simulation environments, overlooking the accuracy of model mimicing step-by-step human behavior. To date, no prior work has focused on objectively evaluating model-generated action traces–that is, assessing whether a model’s sequence of decisions faithfully aligns with human behavior at each step.
@@ 3.1 Task Definition
To ensure the adaptability to unseen websites, we define and implement a simplified HTML format as our context representation. This format removes non-relevant elements such as scripts, CSS, and purely visual components while preserving essential structural information. Using simplified HTML offers several advantages over custom formats or markdown-based representations: (1) important structural elements, such as lists and tables, remain intact, and (2) LLMs are already familiar with the HTML format, eliminating the need to redefine common elements like ‘button’ and ‘input’.
@@
To comprehensively evaluate the performance of our model, we also evaluate the final outcomes of the shopping session. The final outcomes consist of either a purchase action or a termination action. Since the final outcome is always the last action in a session, it marks the conclusion of the user’s shopping journey. This evaluation setup allows for a more detailed analysis of how accurately the model can generate both process-centric action-level behaviors and the final shopping outcomes.
@@ 3.2 Synthesized Reasoning Trace
Reasoning traces are crucial for understanding users’ action choices but are difficult to collect; thus, they are often not available in behavioral datasets. We employ a reasoning synthesis pipeline to generate them using an LLM. To guide the reasoning generation process, we provide the LLM with the observation context and the corresponding action. Additionally, we record a real human customer’s think-aloud shopping sessions as in-context learning examples.
The LLM is then prompted to generate a free-text reasoning explaining the user’s decision. This approach ensures that the synthetic reasoning traces are both coherent with the action and reflective of human reasoning. By bridging the gap between raw behavioral data and the reasoning processes, our method enhances the explainability of the model.
@@ 3.3 Model Architecture
To incorporate these enriched action traces, we build on existing pre-trained LLMs as our base models. The input to the model consists of two components: (1) a sequence of historical contexts (what the user observed), and (2) the corresponding actions and generated reasoning trace (what the user did and why). The model is trained to generate the next reasoning and action, enabling it to simulate human behavior with greater fidelity.
During the training stage, the model receives the full sequence of a user session—including context, synthetic reasoning, and action—as a single concatenated input. The training objective is to minimize the next-token prediction loss for the reasoning and action tokens, while the loss for the context tokens is masked out. Subsequently, in the evaluation stage, the model is provided with historical context and past reasoning traces and actions. The model first generates the reasoning for the next action; then, based on the generated reasoning, it generates the next action.
@@ 4.1 Dataset Construction
Our training dataset was constructed using data from one of the largest e-commerce platforms globally. The dataset contains 31,865 user sessions from 3,526 users in the online shopping scenario, comprising 230,965 user actions. The session’s final outcome includes 4,432 purchase actions and 27,433 session termination actions. We leveraged our data synthesis pipeline detailed in Section 3.2 to generate the synthetic reasoning for each action based on the context using Claude-3.5-Sonnet.
@@ 4.2 Evaluation and Metrics
This setup allows us to evaluate the model’s performance in a sequential reasoning setting, closely mirroring how real users behave during the session. This process resulted in a total of 357 sessions with 932 test cases.
@@
To compare against our fine-tuned models, we evaluated a set of strong instruction-tuned LLMs under the in-context learning (ICL) setting (a.k.a., prompt-based setting). Specifically, we used several variants of Claude, LLaMA, and Mistral as representatives for general-purpose pre-trained LLMs, along with DeepSeek-R1 as a representative for reasoning LLMs.
@@
colspec = X[l]cc,
cells=m,
cell10,12,141=r,font=,
[…]
(Acc.) Outcome
[…]
Claude 3.5 Sonnet 9.72% 15.91%
Claude 3.7 Sonnet 9.34% 12.81%
DeepSeek-R1 11.86% 20.01%
Qwen2.5-7B 4.25% 11.94%
Mistral-7B-v0.3 4.25% 11.27%
Llama 3.2 3B 2.93% 8.60%
[…]
w/o reasoning 16.67% 26.92%
Mistral-7Bv0.3 SFT 15.84% 30.12%
w/o reasoning 14.17% 17.99%
Llama-3.2-3B SFT 15.77% 33.99%
w/o reasoning 9.31% 4.73%
[…]
@@
For Next Action Generation, we evaluate whether the model can accurately generate user actions during the process. We use an exact match accuracy, where a generation is considered correct only if the action type, action target (e.g., search box or product link), and action attribute (if applicable, such as search keyword) exactly match the ground truth. To avoid skewing the results toward longer sessions, we compute per-session accuracy first and then average across sessions, ensuring each session contributes equally to the final score.
@@ 4.3 Experimental Setup
Model training was performed on a GPU cluster consisting of NVIDIA H200 GPUs, with each training job utilizing eight nodes, each equipped with eight GPUs, for a total of 64 GPUs, each with 140 GB of GPU memory.
@@ 4.4 Evaluation Results and Analysis
Following previous works Lutz et al. (2024); Deng et al. (2023), we evaluate the performance of prompt-based LLMs in simulating human behavior on shopping scenarios, with results presented in Table 1. Our findings indicate that while state-of-the-art LLMs have demonstrated strong capabilities in various tasks and domains, their ability to accurately simulate human behaviors remains limited. Compared to general instruction-tuned models, reasoning-focused DeepSeek-R1 achieved a higher accuracy of 11.86% in the action generation task, suggesting that incorporating reasoning mechanisms offers some advantage. Similarly, for final outcome prediction, DeepSeek-R1 attained an F1 score of 20.01%, outperforming other baseline models. These results highlight the challenges of human behavior simulation and suggest that reasoning models may provide incremental improvements, though significant gaps still remain.
We then fine-tuned LLaMA 3.2, Qwen 2.5, and Mistral using our training dataset. The results demonstrate that fine-tuning LLMs with action traces and synthesized reasoning traces significantly enhances performance. Qwen 2.5-7B achieved 17.26% accuracy in action generation, significantly surpassing DeepSeek-R1 by 5.4% (, McNemar’s test). Similarly, LLaMA 3.2-3B reached an F1 score of 33.99% on the final outcome prediction task, further confirming the effectiveness of fine-tuning. All fine-tuned models significantly outperformed their own ICL variants (, McNemar’s test). These findings underscore that incorporating domain-specific fine-tuning with synthesized reasoning traces leads to substantial improvements in the accuracy of human behavior simulation.
We also visualize the distribution of user actions across different models compared to real human behavior in Figure 2. The fine-tuned models exhibit a distribution that more closely aligns with human action patterns, capturing a more natural balance between search, product clicks, and filter actions. In contrast, Claude 3.5 Sonnet and DeepSeek-R1 displays a noticeably higher purchase rate while conducting only a single search per session, diverging from actual user behavior. This misalignment suggests that fine-tuning not only improves accuracy but also helps models better replicate realistic shopping behaviors rather than over-prioritizing purchase actions.

### v7

# Can LLM Agents Simulate Multi-Turn Human Behavior? Evidence from Real Online Customer Behavior Data

@@ Abstract
Recent research shows that LLM Agents can generate “believable” human behaviors via prompt-only methods, and such agents have been increasingly adopted in downstream applications. However, existing evaluation of these agents only focuses on qualitative believability (whether human raters think they are accurate), leaving open questions of whether LLM agents can accurately generate step-by-step actions mimicking a particular human’s behavior in a multi-turn interaction task. In this work, we take shopping as a case study and present the first large-scale quantitative evaluation of state-of-the-art LLMs’ ability to accurately simulate human behavior. Using real-world data from 31,865 online shopping sessions containing 230,965 user actions, our evaluation reveals that prompt-based LLMs (DeepSeek-R1, Llama, Claude) achieve only 11.86% accuracy in generating human actions, highlighting a substantial gap in actual behavioral accuracy. Through experiments, we also showcase that strategies as simple as fine-tuning LLMs on real human click-through data augmented with synthesized reasoning traces can greatly enhance models’ performance. The fine-tuned Qwen2.5-7B achieves 17.26% action generation accuracy and 33.86% F1 score on final purchase prediction, representing substantial improvements of 5.4% and 13.85% over prompt-only baselines. This work establishes the first rigorous benchmark for human behavior simulation and provides actionable insights for developing more accurate LLM agents for future downstream applications.
ode]html
Can LLM Agents Simulate Multi-Turn Human Behavior? Evidence from Real Online Customer Behavior Data
@@ 1 Introduction
Recent advances in large language models (LLMs) have enabled the simulation of human behavior across a range of applications, including web automation Gur et al. (2023); Zhou et al. (2024), social interaction behaviors Park et al. (2023), interpersonal trust behaviors Xie et al. (2024), and user interface interactions Taeb et al. (2024). These developments have sparked a growing interest in developing LLM Agent, agentic AI systems that uses LLMs to simulate human behaviors Chen et al. (2025a), in many domains, such as social science simulation Park et al. (2024), UI/UX testing Lu et al. (2025a), and expert annotation simulation Chen et al. (2025b). Specifically, in the online shopping domain, researchers have used LLM Agents as virtual customers to test website features Lu et al. (2025a), conduct automated A/B testing on design variants Wang et al. (2025), and evaluate agentic AI systems Sun et al. (2025). These applications build on the finding that LLM Agents are able to simulate “believable” human behavior Park et al. (2023, 2024), whether those behaviors are aligned with step-by-step human actions remains unclear. Thus, a fundamental question remains unanswered: How accurate can LLMs truly replicate human behavior?
@@
In this paper, we focus specifically on the human behavior simulation task that is to generate the next action the user is most likely to perform in a multi-turn interaction session, based on the current observation and the history of past actions. For instance, in an online shopping scenario, the model observes the current webpage context (e.g., a product list) and the user’s action history (e.g., previous clicks or queries), and generates the next plausible action a human would take (e.g., add some product to the shopping cart).
Existing evaluations Park et al. (2023) of human behavior simulation primarily emphasize subjective measures of “believability” (“how much people feel it is like a human”) rather than the objective “accuracy” (“how much it acts like a human”). The most relevant works that measure the objective model accuracy focus only on the final outcome of a task (e.g., purchasing the final product or not Yao et al. (2023), or ultimately trusting the partner or not Xie et al. (2024)), without examining whether the intermediate decision and action sequences align with those of actual humans. Consequently, the field currently lacks a robust and quantitative understanding for assessing LLMs at the process-centric, action-level simulation of human behaviors.
@@
To bridge this gap, we take online shopping as a case study and provide the first systematic evaluation of SOTA LLMs’ accuracy in process-centric, action-level behavior simulation tasks. We leverage a large-scale, real-world dataset consisting of 31,865 user click-through sessions from 3,526 users on an online shopping platform111Platform name omitted for anonymity. Each shopping session (Figure 1) comprises a series of timestamp-aligned context, actionpairs, where the context reflects the webpage observed by the user (e.g., product views, filter states), and the action denotes user inputs such as clicks, searches, or session termination actions. In total, the dataset has 230,965 user actions, and the final outcomes of the sessions include 4,432 purchase actions and 27,433 session termination actions. This dataset enables us to rigorously evaluate how accurately various LLMs can generate human-like behaviors at the action level. We evaluate different models on the next action generation task, benchmarking both the accuracy of generated actions throughout the session and the F1 score for final session outcome prediction (i.e. purchase or not), following protocols similar to existing work on modeling online shopping behavior.
@@
Beyond evaluation, our dataset uniquely positions us to fine-tune LLMs to enhance their accuracy in behavior simulation tasks. While prior work has primarily relied on prompt-based approaches, we show that simply fine-tuning the model with user click-through data leads to significantly better accuracy in action generation and session outcome prediction. Furthermore, drawing inspiration from reasoning-augmented modeling DeepSeek-AI et al. (2025), we hypothesize that exposing models to intermediate reasoning traces–even if synthetically generated–can enhance their ability to simulate human behavior. To test this hypothesis, we augment our dataset with synthesized reasonings from action traces using Claude 3.5 Sonnet and fine-tune models using this augmented data (context, action, reasoning triplets) to learn how to generate not only accurate actions but also the underlying reasoning. Our results show that this reasoning-augmented fine-tuning further boosts model performance, highlighting the importance of modeling not just what humans do, but also why they do it. Taken together, our results provide the first quantitative evidence that out-of-the-box LLMs cannot accurately simulate action-level human behavior in realistic settings, and our ablation and error analyses provide insights to support future research.
In summary, this paper has three main contributions:
@@
We conduct the first quantitative and process-centric evaluation of LLMs’ ability to simulate human web action behaviors using real-world online shopping data.
We show that out-of-the-box LLMs cannot accurately predict human behavior in the next action prediction setting, while simple fine-tuning achieves substantially better performance.
We demonstrate that fine-tuning LLMs with synthesized reasoning traces further enhances their accuracy, highlighting the importance of modeling not only what humans do but also why they do it for faithful human behavior simulation.
@@ 2.1 Simulation of Human Behavior with LLM
However, the evaluation of these works remains limited in scope. Some focus on the subjective believability of process-centric action traces. For instance, Lu et al. (2025a) conducted qualitative interviews to assess participants’ perceptions of the realism of their UXAgent system. Similarly, Park et al. (2023) proposed an evaluation framework that identified emergent social behaviors among generative agents. On the other hand, works that pursue objective evaluation often do so in a single-shot, outcome-centric manner. Zhou et al. (2024) introduced WebArena, a controlled environment for benchmarking web agents based on task completion rates. ReAct Yao et al. (2023) measured success rates in simulation environments, overlooking the accuracy of model mimicing step-by-step human behavior. To date, no prior work has focused on objectively evaluating model-generated step-by-step actions–that is, assessing whether a model’s sequence of decisions faithfully aligns with human behavior at each step.
@@ 3.1 Task Definition
To ensure the adaptability to unseen websites, we define and implement a simplified HTML format as our context representation. This format removes non-relevant elements such as scripts, CSS, and purely visual components while preserving essential structural information. Using simplified HTML offers several advantages compared to custom defined formats (domain specific language, DSLs) or markdown-based representations: (1) important structural elements, such as lists and tables, remain intact, and (2) LLMs are already familiar with the HTML format, eliminating the need to redefine common elements like “button” and “input”.
@@ 3.2 Synthesized Reasoning Trace
Reasoning traces are crucial for understanding users’ action choices but are difficult to collect; thus, they are often not available in behavioral datasets. We employ a reasoning synthesis pipeline to generate them using an LLM. To guide the reasoning generation process, we provide the LLM with the observation context and the corresponding action. Additionally, we record real human customers’ think-aloud shopping sessions Eccles and Arsal (2017) as in-context learning examples. We then prompt the LLM to generate a free-text reasoning explaining the user’s decision. Following recent works on reasoning-augmented LLMs Wei et al. (2023); DeepSeek-AI et al. (2025), the synthesized reasoning is not intended to replicate the actual human thought process. Instead, its purpose is to enhance the model’s predictive accuracy by providing structured intermediate representations that help the LLM better align actions with contextual cues. This approach ensures that the reasoning traces are coherent with the observed actions while improving the model’s behavioral fidelity and explainability.
@@ 3.3 Model Architecture
To incorporate these enriched action traces, we build on existing pre-trained LLMs as our base models. The input to the model consists of two components: (1) a sequence of historical contexts (what the user observed), and (2) the corresponding actions and generated reasoning trace (what the user did and why). The model is trained to generate the next reasoning corresponding action, enabling it to simulate human behavior with greater fidelity.
During the training stage, the model receives the full sequence of a user session—including context, synthetic reasoning, and action—as a single concatenated input. The training objective is to minimize the next-token prediction loss for the reasoning and action tokens, while the loss for the context tokens is masked out. Subsequently, in the evaluation stage, the model works in an autoregressive manner: it is provided with historical context and past reasoning traces and actions, and is asked to first generate the reasoning for the next action; then, based on the generated reasoning, it generates the next action.
@@ 4.1 Dataset Construction
Our training dataset was constructed using data from one of the largest e-commerce platforms globally 222Platform name omitted for anonymity. The dataset contains 31,865 user sessions from 3,526 users in the online shopping scenario, comprising 230,965 user actions. The session’s final outcome includes 4,432 purchase actions and 27,433 session termination actions. We leveraged our data synthesis pipeline detailed in Section 3.2 to generate the synthetic reasoning for each action based on the context using Claude-3.5-Sonnet. The dataset was derived from traffic logs of a small group of users who explicitly opted into a beta testing feature and consented to data collection during the process. Users who opted out of data collection were excluded from the dataset. The data was anonymized to remove any personally identifiable information during processing. Due to the sensitive nature of the data and to protect user privacy, we are unable to release the dataset.
@@
| Model | Generated Next Action | Session Outcome | ||||
| Accuracy | vs Base | v.s. DS-R1 | F1 Score | vs Base | v.s. DS-R1 | |
| Open-Source Models | ||||||
| DeepSeek-R1 | 11.86% | - | - | 20.01% | - | - |
| Llama 3.1 8B | 5.05% | - | -6.81% | 10.87% | - | -9.14% |
| Llama 3.1 70B | 8.19% | - | -3.67% | 12.69% | - | -7.32% |
| Mixtral 8x7B | 5.41% | - | -6.45% | 13.16% | - | -6.85% |
| Qwen2.5-7B | 4.25% | - | -7.61% | 11.94% | - | -8.07% |
| Qwen2.5-3B | 3.91% | - | -7.95% | 10.87% | - | -9.14% |
| Qwen2.5-1.5B | 3.27% | - | -8.59% | 7.94% | - | -12.07% |
| Mistral-7B-v0.3 | 4.25% | - | -7.61% | 11.27% | - | -8.74% |
| Llama 3.2 3B | 2.93% | - | -8.93% | 8.60% | - | -11.41% |
| Llama 3.2 1B | 3.71% | - | -8.15% | 3.09% | - | -16.92% |
| Proprietary Models | ||||||
| Claude 3.5 Haiku | 9.18% | - | -2.68% | 14.77% | - | -5.24% |
| Claude 3 Opus | 6.78% | - | -5.08% | 15.08% | - | -4.93% |
| Claude 3 Sonnet | 8.42% | - | -3.44% | 17.40% | - | -2.61% |
| Claude 3.5 Sonnet | 9.72% | - | -2.14% | 15.91% | - | -4.10% |
| Claude 3.5 Sonnet v2 | 11.69% | - | -0.17% | 18.54% | - | -1.47% |
| Claude 3.7 Sonnet | 9.34% | - | -2.52% | 12.81% | - | -7.20% |
| Fine-tuned Models | ||||||
| Qwen2.5-7B | 16.67% | - | 4.81% | 26.92% | - | 6.91% |
| + reasoning | 17.26% | 3.54% | 5.40% | 33.86% | 25.78% | 13.85% |
| Mistral-7B-v0.3 | 14.17% | - | 2.31% | 17.99% | - | -2.02% |
| + reasoning | 15.84% | 11.79% | 3.98% | 30.12% | 67.43% | 10.11% |
| Llama-3.2-3B | 9.31% | - | -2.55% | 4.73% | - | -15.28% |
| + reasoning | 15.77% | 69.39% | 3.91% | 33.99% | 618.60% | 13.98% |
@@ 4.2 Evaluation and Metrics
To assess pre-trained models’ ability in predicting human behavior, we evaluated a set of instruction-tuned LLMs under the in-context learning (ICL) setting (a.k.a., prompt-based setting). Specifically, we used several variants of Claude, LLaMA, and Mistral as representatives for general-purpose pre-trained LLMs, along with DeepSeek-R1 as a representative for reasoning LLMs. To preserve user privacy, we run open-source model on our own GPU cluster and
@@
For Next Action Generation, we evaluate whether the model can accurately generate user actions during the process. We use an exact match accuracy, where a generation is considered correct only if the action type, action target (e.g., search box or product link), and action attribute (such as search keyword) exactly match the ground truth. To avoid skewing the results toward longer sessions, we compute per-session accuracy first and then average across sessions, ensuring each session contributes equally to the final score.
@@ 4.3 Experimental Setup
Model training was performed on a GPU cluster consisting of NVIDIA H200 GPUs, with each training job utilizing eight nodes, each equipped with eight GPUs, for a total of 64 GPUs, each with 140 GB of GPU memory. A typical job takes 3 hours on 64 GPUS, and in total, we used about 3700 H200 GPU hours for our experiments.
@@ 4.4 Evaluation Results and Analysis
Following previous works Lutz et al. (2024); Deng et al. (2023), we evaluate the performance of prompt-based LLMs in simulating human behavior on shopping scenarios, with results presented in Table 1. Our findings indicate that while state-of-the-art LLMs have demonstrated strong capabilities in various tasks and domains, their ability to accurately simulate human behaviors remains limited, with Claude 3.5 sonnet v2 achieving 11.69% accuracy in next action prediction. Compared to general instruction-tuned models, reasoning-focused DeepSeek-R1 achieved a higher accuracy of 11.86% in the action generation task, suggesting that incorporating reasoning mechanisms offers some advantage. Similarly, for final outcome prediction, DeepSeek-R1 achieved an F1 score of 20.01%, outperforming other baseline models. Additionally, larger model usually performs better than their smaller variants. These results indicate that reasoning ability positively impacts performance on human behavior prediction tasks, supporting the hypothesis that such tasks benefit from models trained with reasoning-oriented objectives.
We then fine-tuned various smaller open-sourced LLMs, including LLaMA 3.2, Qwen 2.5, and Mistral, using our training dataset. The results demonstrate that fine-tuning LLMs with action traces and synthesized reasoning traces significantly enhances performance. Qwen 2.5-7B achieved 17.26% accuracy in action generation, significantly surpassing DeepSeek-R1 by 5.4% (, McNemar’s test). Similarly, LLaMA 3.2-3B reached an F1 score of 33.99% on the final outcome prediction task, further confirming the effectiveness of fine-tuning. Additionally, all fine-tuned models significantly outperformed their own ICL variants (, McNemar’s test). These findings underscore that incorporating domain-specific fine-tuning with synthesized reasoning traces leads to substantial improvements in the accuracy of human online shopping behavior simulation.
Figure 2 compares the distribution of actions generated by different models with real human behavior. Human users rarely apply filters and instead rely heavily on iterative search, averaging 2.82 searches per session, which is over seven times more frequent than filter actions. This reflects natural behaviors such as correcting typos and revising keywords (as shown in Sec. 4.6). In contrast, pre-trained LLMs such as Claude 3.5 Sonnet and DeepSeek-R1 tend to stick to the initial search keyword without revision, overuse filtering actions, and produce disproportionately high purchase rates that diverge from actual user behavior. We hypothesize that this bias arises because existing LLM agent benchmarks like WebShopYao et al. (2022) and WebArenaZhou et al. (2024) primarily evaluate task completion (i.e., making a purchase), which incentivizes models to optimize for purchase-heavy trajectories rather than realistic browsing and search behavior.
The fine-tuned models exhibit a distribution that more closely aligns with human action patterns, capturing a more natural balance between search refinement, product clicks, and minimal reliance on filtering. Unlike pre-trained models, they do not prematurely converge on purchase-oriented behavior and instead reflect the exploratory and iterative nature of real user sessions.
| Example 1 | Example 2 | |
| Previous Action | search for “disney gift” | search for “tee conector” |
| Human Next Action | search for “disney gift card” | search for “tee connector” |
| Qwen-2.5-7B | search for “disney gift card” | search for “tee connector” |
[…]